An escalator is 15 m long. If a person stands on the escalator, it takes 50 s to ride to the top.
(a) If a person walks up the moving escalator with a speed of 0.600 m/s relative to the escalator, how long does it take the person to get to the top?

(b) If a person walks down the "up" escalator with the same relative speed as in part (a), how long does it take to reach the bottom?
1
Expert's answer
2020-10-01T10:10:24-0400

From the condition, find the velocity of the escalator:


"v=15/50=0.3\text{ m/s}."

(a) In this case, person's velocity adds to the escalator's one, so it will take a shorter time:


"t_a=\frac{l_\text{escalator}}{v_\text{escalator}+v_\text{person}},
\\\space\\
t_a=15/(0.3+0.6)=16.7\text{ s}."

(b) In this case, use a similar equation:


"t_b=\frac{l_\text{perso}}{v_\text{person}-v_\text{escalator}},
\\\space\\
t_b=15/(0.6-0.3)=50\text{ s}."